OpenStudy (anonymous):

The demand equation for a product is p= (q - 4)² And the supply equation is p= q² + q + 7 Where p (in thousands of dollars) is the price per 100 units when q hundred units are demanded and supplied. Determine consumer’s surplus and producer’s surplus under market equilibrium.

OpenStudy (gorv):

q^2+q+7=q^2+16-8q 9q=9 q=1 p=9